OpenStudy (anonymous):

When \frac{8668}{25}+\frac{4141}{9}-\frac{5533}{25} is computed and written as a mixed number in simplest form, what is the fractional part of that mixed number?

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (skullpatrol):

$$\Huge \frac{8668}{25}+\frac{4141}{9}-\frac{5533}{25} $$

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so

OpenStudy (skullpatrol):

What do you think you need to do first?

OpenStudy (anonymous):

make all of the denominators the same

OpenStudy (skullpatrol):

Yes, but look two are already the same, right? :)

OpenStudy (anonymous):

yeah

OpenStudy (skullpatrol):

so combine them into one

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

how do you combine them?

OpenStudy (skullpatrol):

$$\Huge \frac{a}{c} - \frac{b}{c} = \frac{a-b}{c}$$

OpenStudy (anonymous):

then

OpenStudy (skullpatrol):

what grade is this for?

OpenStudy (anonymous):

6th grade

OpenStudy (skullpatrol):

what is $$\Huge \frac{8668}{25}-\frac{5533}{25} $$

OpenStudy (skullpatrol):

$$\Huge \frac{8668-5533}{25} $$

OpenStudy (anonymous):

627/5

OpenStudy (skullpatrol):

VERY good :D

OpenStudy (anonymous):

ok:)

OpenStudy (skullpatrol):

now you need to add that to 4141/9

OpenStudy (skullpatrol):

As you said earlier you need a common denominator

OpenStudy (anonymous):

45

OpenStudy (skullpatrol):

yep

OpenStudy (anonymous):

every time I do the question I get a different answer

OpenStudy (skullpatrol):

so that means you keep making a mistake each time in the arithmetic

OpenStudy (anonymous):

let me try again

OpenStudy (anonymous):

I keep getting 26348/45 but it says it's incorrect

OpenStudy (anonymous):

help!!

OpenStudy (skullpatrol):

What does the question ask for?

OpenStudy (anonymous):

to add 627/5 and 4141/9

OpenStudy (skullpatrol):

is computed and written as a mixed number in simplest form, what is the fractional part of that mixed number?

OpenStudy (anonymous):

oh I totally forgot that part

OpenStudy (anonymous):

is it 23/45?

OpenStudy (anonymous):

are you still there?

OpenStudy (skullpatrol):

Correct :D